# Market Dynamics, Governance and Open Research Metadata in the AI Era

### 21 Apr 2026 v1

@@ I Introduction
These two features interact to produce what we call the innovation annulus [29]—a zone between the open core of structured data that is either free or indistinguishable from free for the user, and the advancing frontier of commercially refined knowledge products (see Fig. 1). As open access becomes more the dominant mode of publication, the annulus cannot be solely positioned as a pathology nor as an information asymmetry imposed by rent-seeking commercial actors. Rather, it is a structural consequence of production friction in a system where the scholarly record was never designed for structured data consumption, and where the definition of what constitutes useful structured data is continuously evolving. The width of the annulus is, in a precise sense, a measure of the system’s distance from perfect efficiency—and because perfect efficiency is logically impossible in a system where useful data types evolve faster than the production system can standardise them, the annulus is a permanent feature of the landscape. In the AI era in which we now sit, we might expect AI to reduce production friction, but given the complexity of the legal rights surrounding text and data-mining, and the challenges of understanding the algorithmic provenance of a piece of data so that it can be trusted [57], it is not clear that this assumption is well-founded.
@@ II Production Friction and the Efficient Market Analogy
The production of structured scholarly data is often discussed as though it were a problem that technology has essentially solved, or will shortly solve, leaving only political and institutional barriers to full openness. This view underestimates the depth and persistence of the frictions involved.
@@ III.1 Competitive strategy and information asymmetry
Research institutions compete for funding, talent and reputation. This competition is not a market distortion but a deliberate feature of research systems designed to allocate scarce resources to the most productive groups [16]. Competition creates demand for information asymmetries: institutions that can identify emerging research strengths more rapidly than and prior to their competitors, benchmark their performance more precisely, or anticipate funder priorities more accurately gain a structural advantage. This mechanism leads to the creation of centres of excellence and a varied landscape with specialisms. This dynamic is analogous to the knowledge spillovers that Jaffe [32] demonstrated in industrial R&D: structured data about who is doing what research, and where, is the mechanism through which competitive intelligence flows—and actors who can access and interpret that data more effectively gain a systematic advantage.
@@ III.4 Reading the annulus geometry
These two properties interact to create four characteristic configurations.
@@ IV.1 A welfare framework
Cost of structuring. Let $C(r_{o})$ denote the full cost of producing and sustaining structure up to level $r_{o}$—entity resolution, disambiguation, classification, and the ongoing stewardship required to keep the structured representation current as the underlying record evolves [18, 10]. $C$ is increasing and convex: the easy structuring tasks (for example, applying DOIs, resolving common institutional names) are accomplished first, and each additional unit of frontier structuring requires more domain expertise, more validation and more continuing maintenance. We do not separate production from maintenance of structure because they are incurred jointly and scale together with $r_{o}$.
Openness premium. Let $B(r_{i})$ denote the additional welfare gain from making the range $0,r_{i}]$ openly available rather than providing it on restricted terms—the equity, access and knowledge-spillover premium that openness confers on top of the base structuring value already captured in $V$. $B$ is increasing and concave in $r_{i}$: the first units of open provision (basic bibliographic metadata, core identifiers) have enormous marginal value because they set a floor for all institutions, while later units have diminishing marginal impact. This is where the equity argument of Sec. III.5 enters formally—the social weight on the early units of $B$ is high because they serve under-resourced institutions that would otherwise be excluded [61].
@@ IV.3 Relation to patent theory and limitations
We note these limitations not to undermine the framework but to identify the empirical and theoretical work required to make it operational. Even in its current form, the framework provides something that the governance conversation around open research information has lacked: a formal language for reasoning about the trade-offs involved in setting the boundary between open and commercial, and a set of testable predictions about how the annulus should respond to changes in technology, demand and policy.
@@ V.1 Frontier acceleration and cost reduction
AI techniques—automated metadata extraction, entity resolution, semantic classification, citation parsing—have dramatically lowered the cost and increased the speed of basic data structuring. Tasks that previously required large teams of manual curators can now be accomplished computationally at a fraction of the cost. This is unambiguously positive: it pushes the inner boundary of the annulus outward, expanding the volume of data that can be provided openly, so long as it is derived with “appropriate provenance”.
@@ V.2 Quality threshold elevation
A less-discussed effect of AI is that it raises the quality threshold at which data refinement has commercial value. As basic structuring becomes commoditised through AI automation, the annulus does not disappear—it migrates to higher-order refinement tasks. The commercially valuable frontier moves from “can you structure this data at all?” to “can you structure it at a quality level that supports investment decisions, regulatory submissions or competitive strategy?” The bar rises because the floor has risen.
@@ VI Structured In-Paper Metadata: A Frontier Case Study
Funding acknowledgements. Information about who funded a piece of research is present in the vast majority of papers but is structured in wildly inconsistent ways. Some authors include full funder names and grant numbers; others use ambiguous abbreviations; still others include only a brief narrative acknowledgement. Crossref’s funder registry has begun to standardise this, and publishers increasingly request structured funding information—but the legacy record and the inconsistency of current practice mean that producing reliable, analysis-ready funding data at scale requires significant investment in natural language processing and entity resolution.
CRediT author contribution statements. The CRediT taxonomy [6, 12, 46] provides a standardised vocabulary for describing author contributions (conceptualisation, methodology, writing, supervision, etc.). Initiated as a collaboration between the Wellcome Trust and Digital Science, CRediT was designed to address the inadequacy of ordered author lists as a mechanism for attribution and credit [5]. Adoption is growing but uneven. A recent retrospective analysis found that as of 2024, only 22.5% of original research articles with available full text in Dimensions included CRediT role information, with significant variation across publishers, disciplines and countries [4]. Where CRediT statements are present, they are not always machine-readable; where they are machine-readable, they are not always deposited in Crossref metadata. The Dimensions team, for example, has invested in the creation and curation of AI models that identify and extract author contribution statements across the literature—work that operates at accuracy levels that still require improvement and hence further investment [29]. These data would be of significant value to the evaluation community and to anyone involved in tenure and promotion processes, but no universally accepted structured data format yet makes them widely available.
Data availability statements. Many journals now require authors to declare whether and where their research data are available. These statements are typically free-text, with no standard vocabulary or structure. Extracting structured information about data availability at scale—particularly distinguishing between “data available on request,” “data deposited in [specific repository],” and “no data were generated”—requires significant processing.
@@
Each of these data types follows a recognisable lifecycle. They begin as unstructured free text within papers (wide annulus, high production cost for anyone wanting to use them at scale). They move through a standardisation phase in which a taxonomy or identifier system is developed (CRediT, Crossref funder registry). They pass through a phase of uneven adoption in which the standard exists but is not universally applied. And eventually—though this has not yet happened for most of these data types—they become part of the baseline structured record that publishers produce natively, at which point the annulus for that specific data type shrinks toward zero.
@@ VII Governance: Managing the Boundary
If the annulus is a permanent structural feature of the knowledge ecosystem rather than a pathology to be eliminated, then the governance question changes. It is no longer “how do we make everything open?” but “how do we ensure the annulus is the right width?” Too thin, and there is insufficient economic incentive for the frontier data production that serves differentiated demand and drives innovation in data quality. Too thick, and the research ecosystem is poorly served—locked into paying for data whose production cost has fallen below what the price reflects.
@@ VII.1 The Entrepreneurial State and its limits
But applying the Entrepreneurial State model to the full spectrum of data refinement needs creates difficulties that the annulus framework makes visible. The welfare framework in Sec. IV makes the trade-off explicit: the Entrepreneurial State model corresponds to the limiting case in which public subsidy $S$ fully replaces annulus revenue and the annulus width $w$ approaches zero. This is logically coherent but practically demanding for three reasons.
@@ VII.3 The Barcelona Declaration as a norm-setting forum
The Barcelona Declaration on Open Research Information [9, 38] introduces a governance logic that is compatible with the functional annulus model. Rather than mandating specific data to be free, the Declaration frames openness as a civic obligation—what Porter [56, 57] has called “research information citizenship”—and distributes responsibility across producers, consumers and aggregators of metadata. Waltman [67] has argued that responsible research assessment requires open scholarly metadata—a position that the annulus framework refines: the question is not whether metadata should be open in principle, but which metadata types should be inside the open core at any given stage of the system’s maturation.
This framing is significant because it positions the Declaration not as a regulation but as a norm-setting body. In the same way that the scholarly community has developed informal norms about PhD embargoes and data sharing, the Barcelona Declaration could be the venue where the community develops norms about which data types should be inside the open core, what quality standards apply to open metadata, and what responsibilities different actors have in the production chain. Barcelona brings together funders, institutions, and infrastructure providers—arguably the right constituency to develop these norms, though one could ask whether its governance should expand to include the corporate research users whose demand for refined data shapes part of the annulus.
@@ VIII Dimensions: An Illustration of Annulus Dynamics
This experience illustrates several of the paper’s theoretical claims. First, that the annulus has sectoral dimensions: different data types and different quality levels occupy different positions, and the inner boundary moves outward at different rates for different user communities. In the geometric terms of Sec. III.4, Dimensions’ trajectory represents a case where the inner arc has been deliberately pushed further out than market forces alone would produce—a strategic choice to increase the openness ratio for the academic segment of the annulus. Second, that the annulus is shaped by technological capability: as AI-enabled structuring lowers the cost of basic normalisation, data types that were previously commercially valuable become candidates for open provision. Third, that the boundary between open and commercial is not static but is determined dynamically by the interaction of production cost and user demand—and that a commercial actor can, by strategic choice, accelerate the outward movement of the inner boundary. Fourth, that the constraint identified in Sec. III.4—that the inner boundary cannot move outward faster than the outer boundary—operates in practice: the open layers of Dimensions are sustained, in part, by the investment in frontier refinement that drives the outer boundary forward.
The limitations of this model should also be noted. The governance of the boundary between open and commercial layers is determined by the commercial actor rather than by community oversight. The open provision is contingent on the commercial operation remaining viable. And the reliance on a single commercial actor for a significant portion of open data infrastructure creates concentration risks. These limitations reinforce the paper’s argument that governance norms—of the kind the Barcelona Declaration is beginning to develop—are needed to ensure that the boundary is managed in the public interest and not solely according to commercial logic.
@@ IX Open Questions
Optimal annulus width. The welfare framework in Sec. IV yields qualitative predictions about how the optimal annulus width should vary by data type, but making these predictions quantitative requires empirical estimates of the functional forms: the social benefit function $B(r_{i})$, the production cost function $C(r_{o})$, the maintenance cost $M(r_{i})$, and the revenue function $R(w)$. None of these is currently known for any data type in the scholarly ecosystem, and their estimation presents significant methodological challenges—not least because the “radial” dimension of the annulus is not directly measured in any existing dataset. Developing proxy measures—for example, using the time lag between commercial and open availability of specific metadata types as a proxy for annulus width, or using infrastructure provider cost data to estimate MM—would ground the framework quantitatively. The dynamic extension, casting the problem as an optimal control problem with technological change driving the evolution of both boundaries, would connect to the broader literature on optimal technology diffusion and would yield predictions about the trajectory of annulus compression that could be tested against longitudinal data.
@@
Governance machinery for the inner boundary. What institutional mechanisms are required to ensure the inner boundary moves outward at an appropriate rate? The Crossref mechanism works for data types already in the publishing workflow. The Barcelona Declaration provides a normative framework. But Ostrom’s [51] design principles predict that governance without effective monitoring and graduated sanctions will erode under pressure. Empirical study of signatory behaviour change—analogous to the compliance studies conducted for DORA [69]—would test the Declaration’s effectiveness.
@@ X Conclusion
While we have principally confined our attention to the question of scholarly metadata, this model can also be used to consider open access in general and is also well positioned to be a helpful model when thinking about the fundamental changes that are taking place in the software and open source ecosystem where, with the rise of vibe coding, the following comments on AI are particularly pertinent.
AI has changed the parameters of this system profoundly. It has lowered the cost of basic structuring, pushing the inner boundary outward and expanding the volume of data that can be provided openly. But it has also raised the quality threshold at which refinement has value, shifted the commercially relevant frontier to higher-order tasks, and introduced systemic risks through unprovenanced AI-derived metadata. The annulus persists—not because commercial actors universally resist openness, but because the economics of data refinement in a system with technological frictions and differentiated demand make it a structural feature of the landscape.
@@
The annulus will persist. The question is whether we govern it wisely: ensuring that the inner boundary moves outward as technology matures, that the open core is built on principled rather than performative foundations, and that the scholarly data ecosystem serves all its users—including those who cannot afford to pay for frontier refinement—as equitably and efficiently as the state of the art allows.

### 24 Apr 2026 v2

@@ I Introduction
These two features interact to produce what we call the innovation annulus [29]—a zone between the open core of structured data that is either free or indistinguishable from free for the user, and the advancing frontier of commercially refined knowledge products (see Fig. 1). As open access becomes more the dominant mode of publication, the annulus cannot be solely positioned as a pathology nor as an information asymmetry imposed by rent-seeking commercial actors. Rather, it is a structural consequence of production friction in a system where the scholarly record was never designed for structured data consumption, and where the definition of what constitutes useful structured data is continuously evolving. The width of the annulus is, in a precise sense, a measure of the system’s distance from perfect efficiency—and because perfect efficiency is logically impossible in a system where useful data types evolve faster than the production system can standardise them, the annulus is a permanent feature of the landscape. In the AI era in which we now sit, we might expect AI to reduce production friction, but given the copyright complexities of text and data-mining and the challenges of understanding the algorithmic provenance of a piece of data so that it can be trusted [57], it is not clear that this assumption is well-founded.
@@ II Production Friction and the Efficient Market Analogy
The production of structured scholarly data is often discussed as though it were a problem that technology has essentially solved, leaving only political and institutional barriers to full openness. This view underestimates the depth and persistence of the frictions involved.
@@
A deeper physical grounding for this claim is available and worth making explicit. In The Human Use of Human Beings, Wiener framed organised systems—biological, mechanical, institutional, informational—as local enclaves that persist in defiance of the second law of thermodynamics only by continuously importing energy and exporting disorder to their surroundings [68]. Structured information is, in this view, a form of negentropy; any system that maintains it must continuously do work against the universal tendency toward disorder. Left to themselves, messages accumulate noise, classifications decay, identifiers decouple from the entities they name, and standards fragment as they are locally reinterpreted. The three categories of production friction identified above are specific expressions of this tendency rather than peculiarities of the scholarly ecosystem: source heterogeneity is entropy in the generative process; quality decay is entropy accumulating in a previously ordered store; and the continuous emergence of frontier data types reflects the expansion of the informational phase space faster than any one-time investment can contain. The annulus, from this perspective, is the thermodynamic signature of the work being done against that disorder. Its width for any given data type is, in effect, a measure of the energetic cost of holding that region of the scholarly record in an analysable state; eliminating it would require an impossible condition—an information system in which organisation required no ongoing input. Maxwell’s demon cannot get something for nothing, and neither can a metadata infrastructure: interventions that attempt to collapse the annulus do not remove the cost of maintaining order, they only reassign where that cost is paid. The annulus is the price of sustaining an anti-entropic enclave in an informational system that would, without continuous investment, drift toward unstructured noise.
@@ III.1 Competitive strategy and information asymmetry
Research institutions compete for funding, talent and reputation. This competition is not a market distortion but a deliberate feature of research systems designed to allocate scarce resources to the most productive groups [16]. Competition creates demand for information asymmetries: institutions that can identify emerging research strengths more rapidly than and prior to their competitors, benchmark their performance more precisely, or anticipate funder priorities more accurately gain a structural advantage. This mechanism leads to the creation of centres of excellence and a varied landscape with specialisms. This dynamic is analogous to the knowledge spillovers that Jaffe [32] demonstrated in industrial R&D: structured data about who is doing what research, is the mechanism through which competitive intelligence flows, and actors who can access and interpret that data more effectively gain a systematic advantage.
@@ III.4 Reading the annulus geometry
These two properties interact to create four characteristic configurations:
  1. 1.
@@ IV.1 A welfare framework
Cost of structuring. Let $C(r_{o})$ denote the full cost of producing and sustaining structure up to level $r_{o}$—entity resolution, disambiguation, classification, and the ongoing stewardship required to keep the structured representation current as the underlying record evolves [18, 10]. $C$ is increasing and convex: the easy structuring tasks (for example, applying DOIs, resolving common institutional names) are accomplished first, and each additional unit of frontier structuring requires more domain expertise, more validation and continuing maintenance. We do not separate production from maintenance because they are incurred jointly and scale together with $r_{o}$.
Openness premium. Let $B(r_{i})$ denote the additional welfare gain from making the range $0,r_{i}]$ openly available rather than providing it on restricted terms—the equity, access and knowledge-spillover premium that openness is incremental to the value captured in $V$. $B$ is increasing and concave in $r_{i}$: the first units of open provision (basic bibliographic metadata, core identifiers) have enormous marginal value because they set a floor for all institutions, while later units have diminishing marginal impact. This is where the equity argument of Sec. III.5 enters formally—the social weight on the early units of $B$ is high because they serve under-resourced institutions that would otherwise be excluded [61].
@@ IV.3 Relation to patent theory and limitations
We note these limitations not to undermine the framework but to identify the empirical and theoretical work required to make it operational. Even in its current form, the framework provides something that the governance conversation around open research information has lacked: a formal language for reasoning about the trade-offs involved in setting the boundary, and a set of testable predictions about how the annulus should respond to changes in technology, demand and policy.
@@ V.1 Frontier acceleration and cost reduction
AI techniques—automated metadata extraction, entity resolution, semantic classification, citation parsing—have lowered the cost and increased the speed of basic data structuring. Tasks that previously required teams of manual curators can now be accomplished computationally at a fraction of the cost. This is unambiguously positive: it pushes the inner boundary of the annulus outward, expanding the volume of data that can be provided openly, with “appropriate provenance”.
@@ V.2 Quality threshold elevation
A less-discussed effect of AI is that it raises the quality threshold at which data refinement has commercial value. As basic structuring becomes commoditised through AI automation, the annulus does not disappear—it migrates to higher-order refinement tasks. The commercially valuable frontier moves from “can you structure this data at all?” to “can you structure it at a quality level that supports investment decisions, regulatory submissions or competitive strategy?”
@@ VI Structured In-Paper Metadata: A Frontier Case Study
Funding acknowledgements. Information the funding of a piece of research is present in the vast majority of papers but is structured in wildly inconsistent ways. Some authors include full funder names and grant numbers; others use ambiguous abbreviations; still others include only a brief narrative acknowledgement. Crossref’s funder registry has begun to standardise this, and publishers increasingly request structured funding information—but the legacy record and the inconsistency of current practice mean that producing reliable, analysis-ready funding data at scale requires significant investment in natural language processing and entity resolution.
CRediT author contribution statements. The CRediT taxonomy [6, 12, 46] provides a standardised vocabulary for describing author contributions (conceptualisation, methodology, writing, supervision, etc.). CRediT was designed to address the inadequacy of ordered author lists as a mechanism for attribution and credit [5]. Adoption is growing but uneven. A recent retrospective analysis found that as of 2024, only 22.5% of original research articles with available full text in Dimensions included CRediT role information, with significant variation across publishers, disciplines and countries [4]. Where CRediT statements are present, they are not always machine-readable; where they are machine-readable, they are not always deposited in Crossref metadata. The Dimensions team, for example, has invested in the creation and curation of AI models that identify author contribution statements across the literature—work that operates at accuracy levels that still require improvement and hence further investment [29]. These data would be of significant value to the evaluation community and to anyone involved in tenure and promotion processes, but no universally accepted structured data format yet makes them widely available.
Data availability statements. Many journals now require authors to declare whether and where their research data are available. These statements are typically free-text, with no standard vocabulary or structure. Extracting structured information about data availability at scale, particularly distinguishing between “data available on request,” “data deposited in [specific repository],” and “no data were generated” requires significant processing.
@@
Each of these data types follows a recognisable lifecycle. They begin as unstructured free text within papers (wide annulus, high production cost for anyone wanting to use them at scale). They move through a standardisation phase in which a taxonomy or identifier system is developed (CRediT, Crossref funder registry). They pass through a phase of uneven adoption in which the standard exists but is not universally applied. And eventually—though this has not yet happened for most of these data types—they become part of the baseline structured record that publishers produce natively, when the annulus for that data type shrinks toward zero.
@@ VII Governance: Managing the Boundary
If the annulus is a permanent structural feature of the knowledge ecosystem rather than a pathology to be eliminated, then the governance question changes. It is no longer “how do we make everything open?” but “how do we ensure the annulus is the right width?” Too thin, and there is insufficient economic incentive for the frontier data production that serves differentiated demand and drives innovation in data quality. Too thick, and the research ecosystem is poorly served—locked into paying for data above its actual value.
@@ VII.1 The Entrepreneurial State and its limits
But applying the Entrepreneurial State Model (ESM) to the full spectrum of data refinement needs creates difficulties that the annulus framework makes visible. The welfare framework in Sec. IV makes the trade-off explicit: the ESM corresponds to the limiting case in which public subsidy $S$ fully replaces annulus revenue and the annulus width $w$ approaches zero. This is logically coherent but practically demanding for three reasons.
@@ VII.3 The Barcelona Declaration as a norm-setting forum
The Barcelona Declaration [9, 38] introduces a governance logic that is compatible with the functional annulus model. Rather than mandating specific data to be free, the Declaration frames openness as a civic obligation—what Porter [56, 57] has called “research information citizenship”—and distributes responsibility across producers, consumers and aggregators of metadata. Waltman [67] has argued that responsible research assessment requires open scholarly metadata—a position that the annulus framework refines: the question is not whether metadata should be open in principle, but which metadata types should be inside the open core at any given stage of the system’s maturation.
This framing is significant because it positions the Declaration not as a regulation but as a norm-setting body. In the same way that the scholarly community has developed informal norms about PhD embargoes and data sharing, the Barcelona Declaration could be the venue where the community develops norms about which data types should be inside the open core, what quality standards apply to open metadata, and what responsibilities different actors have in the production chain. The Declaration’s membership includes funders, institutions, and infrastructure providers—arguably the right constituency to develop these norms, though one could ask whether it should expand to include the corporate research users whose data needs shape part of the annulus.
@@ VIII Dimensions: An Illustration of Annulus Dynamics
This practical experience illustrates several of this paper’s theoretical claims. First, that the annulus has sectoral dimensions: different data types and different quality levels occupy different positions, and the inner boundary moves outward at different rates for different user communities. In terms of Sec. III.4, Dimensions’ trajectory represents a case where the inner arc has been deliberately pushed further out than market forces alone would produce—a strategic choice to increase the openness ratio for academic metadata in the annulus. Second, that the annulus is shaped by technological capability: as AI-enabled structuring lowers the cost of basic normalisation, data types that were previously commercially valuable become candidates for open provision. Third, that the boundary between open and commercial is not static but is determined dynamically by the interaction of production cost and user demand—and that a commercial actor can, by strategic choice, accelerate the outward movement of the inner boundary. Fourth, that the constraint identified in Sec. III.4—that the inner boundary cannot move outward faster than the outer boundary—operates in practice: the open layers of Dimensions are sustained, in part, by the investment in frontier refinement that drives the outer boundary forward.
The limitations of this model should also be noted. The governance of the boundary between open and commercial layers is determined by the commercial actor rather than by community oversight. Open provision is contingent on the commercial operation remaining viable. And the reliance on a single commercial actor for a significant portion of open data infrastructure creates concentration risks. These limitations reinforce this paper’s argument that governance norms of the kind the Barcelona Declaration is beginning to develop are needed to ensure that the boundary is managed in the public interest and not according to commercial logic.
@@ IX Open Questions
Optimal annulus width. The welfare framework in Sec. IV yields qualitative predictions about how the optimal annulus width should vary by data type, but making these predictions quantitative requires empirical estimates of the functional forms: the social benefit function $B(r_{i})$, the production cost function $C(r_{o})$, the maintenance cost $M(r_{i})$, and the revenue function $R(w)$. None of these is currently known for any data type in the scholarly ecosystem, and their estimation presents significant methodological challenges—not least because the “radial” dimension of the annulus is not directly measured in any existing dataset. Developing proxy measures—for example, using the time lag between commercial and open availability of specific metadata types as a proxy for annulus width, or using infrastructure provider cost data to estimate MM—would ground the framework quantitatively. The dynamic extension, casting the problem as an optimal control problem with technological change driving the evolution of both boundaries, would connect to the literature on optimal technology diffusion and would yield predictions about the trajectory of annulus compression that could be tested using longitudinal data.
@@
Governance mechanism for the inner boundary. What institutional mechanisms could ensure the inner boundary moves outward appropriately? The Crossref mechanism works for data types already in the publishing workflow. The Barcelona Declaration provides a normative framework, but Ostrom’s [51] design principles predict that governance without effective monitoring and graduated sanctions will erode under pressure. Empirical study of signatory behaviour change—analogous to the compliance studies conducted for DORA [70]—would test the Declaration’s effectiveness.
@@ X Conclusion
Although we have principally confined our attention to consideration of scholarly metadata, we believe that this model can also be useful in thinking about the evolution of open access, and the commercial software / open source ecosystem where, with the rise of vibe coding, the following comments on AI are particularly pertinent.
AI has changed the parameters of this system profoundly. It has lowered the cost of basic structuring, pushing the inner boundary outward and expanding the volume of data that can be provided openly. But it has also raised the quality threshold at which refinement has value, shifted the commercially relevant frontier to higher-order tasks, while simultaneously introducing systemic risks through unprovenanced AI-derived metadata. Again, the annulus persists—not through resistance to openness, but because the economics of data refinement in a system with technological frictions and differentiated demand make it a structural feature of the landscape.
@@
The annulus will persist—the width is the price of being able to trust the data to the level needed to meet a given use case. The question is whether we govern it wisely: ensuring that the inner boundary moves outward as technology matures, that the open core is built on principled rather than performative foundations, and that the scholarly data ecosystem serves all its users—including those who cannot afford to pay for frontier refinement—as equitably and efficiently as the state of the art allows.